A random sample of n = 60 measurements is obtained from a population with μ = 192 and σ 43. Describe the sampling distribution for the sample means by computing the μx and σx.


1
Expert's answer
2021-03-30T04:28:40-0400

"μ_x = μ = 192 \\

σ_x = \frac{σ}{\sqrt{n}} = \frac{43}{\sqrt{60}} = 5.55"